For reaction A-->P the intergrated rate is In [A]t = -Kt + In[A]0

(a) What is the order of the reaction?

(b) Draw the graph of In[A]t vs t and show the intercept and slope

(c) Calculate the half life
Answer #1

The complete step wise calculation in detail is given as below:-

@ The given equation is as below: In [A] + = -kt ten (A). O where, at = time, [A] = Initial Concentration From equation , we

since the above equation 0 is a straight lines time (t) In [A]a us a graphy = mx + c- (y vs x where m= slope and Then from equation ① and graph) c= intercept. , we have m-K , slope = m =-k Ce In CAJ.

● From equation , we have If, to ty and [A]; = (A). Then In Cao -(+42) + en C). en [A]. - ln (2) -ln [A] =-k (1/2) on keve) =